# Extended Friction Models for the Physics Simulation of Servo Actuators

### v1

@@ Abstract
Accurate physical simulation is crucial for the development and validation of control algorithms in robotic systems. Recent works in Reinforcement Learning (RL) take notably advantage of extensive simulations to produce efficient robot control. State-of-the-art servo actuator models generally fail at capturing the complex friction dynamics of these systems. This limits the transferability of simulated behaviors to real-world applications. In this work, we present extended friction models that allow to more accurately simulate servo actuator dynamics. We propose a comprehensive analysis of various friction models, present a method for identifying model parameters using recorded trajectories from a pendulum test bench, and demonstrate how these models can be integrated into physics engines.
The proposed friction models are validated on four distinct servo actuators and tested on 2R manipulators, showing significant improvements in accuracy over the standard Coulomb-Viscous model. Our results highlight the importance of considering advanced friction effects in the simulation of servo actuators to enhance the realism and reliability of robotic simulations.
@@ I Introduction
A common approach to ensuring consistency between simulation and reality in robotics is to identify the dynamics of the actuators [4]. Several studies have proposed identifying servo actuator dynamics by recording trajectories on the real system and minimizing the error between simulated and recorded trajectories. This process involves tuning the servo actuator model parameters using methods like iterative coordinate descent [5] or genetic algorithms [6]. While these methods have demonstrated promising results, they are constrained by the complexity of servo actuator dynamics, which can be challenging to accurately capture using simple models.
@@
In this work, we focus on simulating servo actuators using models that accurately capture the effects of friction. Friction is a complex phenomenon that can, in a lubricated environment such as a motor gearbox, be decomposed into two main components: static friction (also known as dry friction) and viscous friction [9]. Static friction is the force that opposes motion when the joint is stationary, caused by the asperities of the surfaces in contact. It represents a threshold force that must be overcome to initiate motion. Once the joint starts moving, static friction decreases, and viscous friction increases in a pre-sliding regime. At higher speeds, viscous friction becomes dominant over static friction, leading to a sliding regime.
Viscous friction is typically modeled as a linear function of velocity, acting as a damper. Static friction, on the other hand, is often modeled as a constant force, a concept originally described by Coulomb in the 18th century [10]. Despite the identification of other effects impacting static friction, such as the Stribeck effect [11] and load dependence [12], the Coulomb-Viscous model remains widely used in popular physics simulators like MuJoCo [2] and Isaac Gym [3]. Consequently, simulations of servo actuators in these environments may not accurately reflect real-world behaviors, particularly when controlled with low gains, which is common in RL applications.
@@
In the following sections, we present the background and motivation for this work, introduce the extended friction models, explain how to simulate servo actuators with these models, and present the results of our experiments. To demonstrate the effectiveness of the proposed models, four distinct servo actuators are identified: Dynamixel MX-64 [13] and MX-106[14], eRob80:50 and eRob80:100, which are both eRob80[15] with respectively 1:50 and 1:100 reduction ratios. We then compare the Coulomb-Viscous model with the extended models on two 2R arms. Such a comparison is presented in Fig. 1, highlighting the importance of accounting for different friction effects in the simulation of servo actuators. Every experiment conducted in this work is presented in the video accompanying this paper [^1] https://youtu.be/P5-Ked8EoWk.
@@ II-B Friction model
Friction can be seen as a force, here a torque, preventing the motion of the system. The most typical model is the Coulomb-Viscous model [11], generally defined as
$\tau_{f}=-K_{v}\dot{\theta}-sign(\dot{\theta})K_{c}$ (2)
Here, $K_{c}$ is the Coulomb friction, that needs to be overcome for the system to start moving. $K_{v}$ is the viscous friction that is countering the velocity and acts as a damper. With this model, $\tau_{f}$ is a passive force countering the direction of the motion. This formulation implies zero friction when $\dot{\theta}=0$, which does not reflect reality. Moreover, the friction becomes discontinuous when the direction of motion changes, which is both physically inaccurate and unsuitable for simulation.
In this work, we define friction as an upper bound. For the Coulomb-Viscous model, we write:
$\mathcal{M}_{1}:\tau_{f}^{m}=K_{v}|\dot{\theta}|+K_{c},$ (3)
where $\tau_{f}^{m}$ is the torque budget available for friction to prevent the motion. In a simulator using time steps of $\Delta t$ seconds, stopping the motion can be interpreted as bringing the velocity to zero at the next time step, and thus having $\dot{\theta}+\ddot{\theta}\Delta t=0$. Based on equation (1), assuming friction is the stopping factor, this would require a torque $\tau_{f,stop}$ of:
$\tau_{f,stop}=-\biggl{(}\frac{J}{\Delta t}\dot{\theta}+\tau_{m}+\tau_{e}\biggr{)}$ (4)
The resistive torque $\tau_{f}$ caused by friction is given by $\tau_{f,stop}$, as computed in equation (4), and limited to the range $[-\tau_{f}^{m},\tau_{f}^{m}]$. The maximum value $\tau_{f}^{m}$ is reached only at the edge of motion or when friction has been overcome and the system is moving.
@@ III Extended friction models
The Coulomb-Viscous model presented in the previous section is typically used in physics simulation. Based on Fig. 3, we can expect this type of model to lack expressiveness to simulate servo actuators. In this section, we describe those effects and provide extended models to account for them.
@@ III-A Stribeck model (5 parameters)
$\mathcal{M}_{2}:\tau_{f}^{m}=K_{v}\dot{\theta}+K_{c}+e^{-\left|\frac{\dot{\theta}}{\dot{\theta}^{s}}\right|^{\alpha}}K_{c}^{s}$ (5)
The parameter $K_{c}^{s}$ is the Stribeck-Coulomb friction, which is an extra static friction when not moving. The Stribeck velocity $\dot{\theta}^{s}$ parametrizes the zone of influence of the Stribeck effect and $\alpha$ its curvature. This model is used with Dynamixel servo actuators in [6].
@@ III-C Stribeck load-dependent model (7 parameters)
$\mathcal{M}_{4}:\tau_{f}^{m}=K_{v}\dot{\theta}+K_{c}+K_{l}|\tau_{m}-\tau_{e}|$ (7)
$+e^{-\left|\frac{\dot{\theta}}{\dot{\theta}^{s}}\right|^{\alpha}}[K_{c}^{s}+K_{l}^{s}|\tau_{m}-\tau_{e}|]$
@@ III-D Directional model (9 parameters)
$\mathcal{M}_{5}:\tau_{f}^{m}=K_{v}\dot{\theta}+K_{c}+|K_{m}\tau_{m}-K_{e}\tau_{e}|$ (8)
$+e^{-\left|\frac{\dot{\theta}}{\dot{\theta}^{s}}\right|^{\alpha}}[K_{c}^{s}+|K_{m}^{s}\tau_{m}-K_{e}^{s}\tau_{e}|]$
@@ III-E Quadratic model (11 parameters)
$\mathcal{M}_{6}:\tau_{f}^{m}=K_{v}\dot{\theta}+K_{c}+|K_{m}\tau_{m}-K_{e}\tau_{e}|$ (9)
$+e^{-\left|\frac{\dot{\theta}}{\dot{\theta}^{s}}\right|^{\alpha}}[K_{c}^{s}+|K_{m}^{s}\tau_{m}-K_{e}^{s}\tau_{e}|+Q]$
@@ IV-D Simulating the pendulum test bench
3 $\tau_{e}\leftarrow mglsin(\theta)$; test bench dynamics
@@ 4 $J\leftarrow ml^{2}+J_{m}$; test bench dynamics
5 $\tau_{f}^{m}\leftarrow\mathcal{M}(\tau_{m},\tau_{e},\dot{\theta}_{k})$; friction model
6 $\tau_{f,stop}\leftarrow-(\frac{J}{\Delta t}\dot{\theta}_{k}+\tau_{m}+\tau_{e})$; torque to stop
7 $\tau_{f}\leftarrow clip(\tau_{f,stop},[-\tau_{f}^{m},\tau_{f}^{m}])$; applied friction
8 $\ddot{\theta}_{k}\leftarrow\frac{\tau_{m}+\tau_{e}+\tau_{f}}{J}$; angular acceleration
@@ V-B Optimization
In order to compute a score for a set of parameters, Algorithm 1 is used to simulate the trajectories previously introduced. The Mean Absolute Error (MAE) between the simulated and measured trajectories is used as a score.
The parameters of the friction and servo actuators models are identified using the CMA-ES genetic algorithm [22], implemented via the optuna Python module [23]. The logs are split into two sets: 75% are used for parameter identification, and 25% are reserved for validation. Each identification process was repeated 3 times, demonstrating consistent convergence to the same score.

### v2

@@ Abstract
Accurate physical simulation is crucial for the development and validation of control algorithms in robotic systems. Recent works in Reinforcement Learning (RL) take notably advantage of extensive simulations to produce efficient robot control. State-of-the-art servo actuator models generally fail at capturing the complex friction dynamics of these systems. This limits the transferability of simulated behaviors to real-world applications. In this work, we present extended friction models that allow to more accurately simulate servo actuator dynamics. We propose a comprehensive analysis of various friction models, present a method for identifying model parameters using recorded trajectories from a pendulum test bench, and demonstrate how these models can be integrated into physics engines. The proposed friction models are validated on four distinct servo actuators and tested on 2R manipulators, showing significant improvements in accuracy over the standard Coulomb-Viscous model. Our results highlight the importance of considering advanced friction effects in the simulation of servo actuators to enhance the realism and reliability of robotic simulations.
@@ I Introduction
A common approach to ensuring consistency between simulation and reality in robotics is to identify the dynamics of the actuators [4]. Several studies have proposed identifying servo actuator dynamics by recording trajectories on the real system and minimizing the error between simulated and recorded trajectories. This process involves tuning the servo actuator model parameters using methods like iterative coordinate descent [5] or genetic algorithms [6][7]. While these methods have demonstrated promising results, they are constrained by the complexity of servo actuator dynamics, which can be challenging to accurately capture using simple models.
@@
In this work, we focus on simulating servo actuators using models that accurately capture the effects of friction. Friction is a complex phenomenon that can, in a lubricated environment such as a motor gearbox, be decomposed into two main components: static friction (also known as dry friction) and viscous friction [10]. Static friction is the force that opposes motion when the joint is stationary. It represents a threshold force that must be overcome to initiate motion. Once the joint starts moving, static friction decreases, and viscous friction increases in a pre-sliding regime. At higher speeds, viscous friction becomes dominant over static friction, leading to a sliding regime.
Viscous friction is typically modeled as a linear function of velocity, acting as a damper. Static friction, on the other hand, is often modeled as a constant force, a concept originally described by Coulomb in the 18th century [11]. Despite the identification of other effects impacting static friction, such as the Stribeck effect [12] and load dependence [13], the Coulomb-Viscous model remains widely used in popular physics simulators like MuJoCo [2] and Isaac Gym [3]. Consequently, simulations of servo actuators in these environments may not accurately reflect real-world behaviors, particularly when controlled with low gains, which is common in RL applications. An enhanced friction model was proposed in [7] to account for some of these effects, but it does not adequately quantify their individual contributions to the simulation improvement.
@@
In the following sections, we present the background and motivation for this work, introduce the extended friction models, explain how to simulate servo actuators with these models, and present the results of our experiments. To demonstrate the effectiveness of the proposed models, four distinct servo actuators are identified: Dynamixel MX-64 [14] and MX-106[15], eRob80:50 and eRob80:100, which are both eRob80[16] with respectively 1:50 and 1:100 reduction ratios. We then compare the Coulomb-Viscous model with the extended models on two 2R arms. Such a comparison is presented in Fig. 1, highlighting the importance of accounting for different friction effects in the simulation of servo actuators. Every experiment conducted in this work is presented in the video accompanying this paper [^1] https://youtu.be/5XPEEKDnQEM.
@@ II-B Friction model
Friction can be seen as a force preventing the motion of the system. In rotational systems, it results in a torque $\tau_{f}$ opposing the direction of the motion. The most commonly used model is the Coulomb-Viscous model [12], generally expressed as
$\tau_{f}=-K_{v}\dot{\theta}-sign(\dot{\theta})K_{c},$ (2)
where $K_{c}$ and $K_{v}$ are two positive parameters representing respectively the Coulomb friction, which must be overcome for the system to start moving, and the viscous friction, which is countering the velocity and acts as a damper. With this model, $\tau_{f}$ is a passive torque countering the direction of the motion. This formulation implies zero friction when $\dot{\theta}=0$, which does not reflect reality. Moreover, the friction becomes discontinuous when the direction of motion changes, which is both physically inaccurate and unsuitable for simulation.
In this work, we define friction as a torque budget $\tau_{f}^{m}$ available to prevent the motion. For the Coulomb-Viscous model, we express it as
$\mathcal{M}_{1}:\tau_{f}^{m}=K_{v}|\dot{\theta}|+K_{c}$ (3)
In a simulator using time steps of $\Delta t$ seconds, stopping the motion can be interpreted as bringing the velocity to zero at the next time step, leading to the condition: $\dot{\theta}+\ddot{\theta}\Delta t=0$.
Based on equation (1), assuming friction is the stopping factor, this would require a torque $\tau_{f}^{stop}$ of:
$\tau_{f}^{stop}=-\biggl{(}\frac{J}{\Delta t}\dot{\theta}+\tau_{m}+\tau_{e}\biggr{)}$ (4)
Thus, the resistive torque $\tau_{f}$ caused by friction is given by $\tau_{f}^{stop}$, as computed in equation (4), and limited to the range $[-\tau_{f}^{m},\tau_{f}^{m}]$:
$\tau_{f}=clip(\tau_{f}^{stop},[-\tau_{f}^{m},\tau_{f}^{m}])$ (5)
@@ III Extended friction models
The Coulomb-Viscous model presented in the previous section is typically used in physics simulation. Based on Fig. 3, we can expect this type of model to lack expressiveness to simulate servo actuators. In this section, we describe those effects and provide extended models to account for them. All parameters introduced in this section are positive.
@@ III-A Stribeck model (5 parameters)
$\mathcal{M}_{2}:\tau_{f}^{m}=K_{v}|\dot{\theta}|+K_{c}+e^{-\left|\frac{\dot{\theta}}{v_{s}}\right|^{\alpha}}K_{c}^{s}$ (6)
The parameter $K_{c}^{s}$ is the Stribeck-Coulomb friction, which is an extra static friction when not moving. The Stribeck velocity $v_{s}$ parametrizes the zone of influence of the Stribeck effect and $\alpha$ its curvature. This model is used with Dynamixel servo actuators in [6].
@@ III-C Stribeck load-dependent model (7 parameters)
$\mathcal{M}_{4}:\tau_{f}^{m}=K_{v}|\dot{\theta}|+K_{c}+K_{l}|\tau_{m}-\tau_{e}|$ (8)
$+e^{-\left|\frac{\dot{\theta}}{v_{s}}\right|^{\alpha}}[K_{c}^{s}+K_{l}^{s}|\tau_{m}-\tau_{e}|]$
@@ III-D Directional model (9 parameters)
$\mathcal{M}_{5}:\tau_{f}^{m}=K_{v}|\dot{\theta}|+K_{c}+|K_{m}\tau_{m}-K_{e}\tau_{e}|$ (9)
$+e^{-\left|\frac{\dot{\theta}}{v_{s}}\right|^{\alpha}}[K_{c}^{s}+|K_{m}^{s}\tau_{m}-K_{e}^{s}\tau_{e}|]$
@@ III-E Quadratic model (11 parameters)
$\mathcal{M}_{6}:\tau_{f}^{m}=K_{v}|\dot{\theta}|+K_{c}+|K_{m}\tau_{m}-K_{e}\tau_{e}|$ (10)
$+e^{-\left|\frac{\dot{\theta}}{v_{s}}\right|^{\alpha}}[K_{c}^{s}+|K_{m}^{s}\tau_{m}-K_{e}^{s}\tau_{e}|+Q]$
@@ IV-D Simulating the pendulum test bench
3 $\tau_{e,k}\leftarrow mglsin(\theta_{k})$; test bench dynamics
@@ 4 $J\leftarrow ml^{2}+J_{m}$; test bench dynamics
5 $\tau_{f,k}^{m}\leftarrow\mathcal{M}(\tau_{m,k},\tau_{e,k},\dot{\theta}_{k})$; friction model
6 $\tau_{f,k}^{stop}\leftarrow-(\frac{J}{\Delta t}\dot{\theta}_{k}+\tau_{m,k}+\tau_{e,k})$; torque to stop
7 $\tau_{f,k}\leftarrow clip(\tau_{f,k}^{stop},[-\tau_{f,k}^{m},\tau_{f,k}^{m}])$; applied friction
8 $\ddot{\theta}_{k}\leftarrow\frac{\tau_{m,k}+\tau_{e,k}+\tau_{f,k}}{J}$; angular acceleration
@@ V-B Optimization
The parameters of the friction and servo actuators models are estimated using the CMA-ES genetic algorithm [23], implemented via the optuna Python module [24]. Algorithm 1 is used to simulate the trajectories previously introduced, and the Mean Absolute Error (MAE) between the simulated and measured trajectories serves as the cost function for the optimization.
The logs are divided into two sets: 75% for parameter identification and 25% for validation. The identification process for each model was repeated three times, consistently converging to the same scores and parameters in under 5 minutes and approximately 4000 iterations.